Mannix Corporation stock currently sells for $80 per share. The market requires a return of 11 percent on the firm's stock. If the company maintains a constant 7 percent growth rate in dividends, what was the most recent dividend per share paid on the stock?

A) $14.47
B) $2.99
C) $2.87
D) $3.20
E) $5.11
Answer #1

Ans B) $2.99

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
80 = D1 / (11%- 7%)
80 = D1 / 4%
D1 = 80 * 4%
D1 = 3.20
D0 = D1 / (1 + g)
D0 = 3.20 / (1 + 7%)
D0 = 2.99